OpenStudy (anonymous):

Solve: 16x2 − 81 = 0. x = 4, 9 x = −4, −9 x = 9 over 4, x = −9 over 4 x = 81 over 16, x = −81 over 16

pooja195 (pooja195):

Add 81 to both sides. 16x^2−81+81=0+81 16x^2=81

OpenStudy (anonymous):

Thank you :D

pooja195 (pooja195):

No lol hold up

OpenStudy (anonymous):

Oh ok

pooja195 (pooja195):

from there you divde both sides by 16 since we are isolating x

pooja195 (pooja195):

You are left with x= 81/16

OpenStudy (anonymous):

Thanks :D

pooja195 (pooja195):

so whats the answer? ;p